OpenStudy (anonymous):

Find the interval on which the curve of y equals the integral from 0 to x of 2 divided by the quantity 1 plus2 times t plus t squared, dt is concave up.

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

https://gyazo.com/4935cc6bed8b8ff7962d6a8c115c5cc7

OpenStudy (yttrium):

I think you can use integration using partial fractions.

OpenStudy (anonymous):

I took the second derivative.

OpenStudy (anonymous):

@Yttrium you could, but the fundamental theorem of calc is better in this situation. (thanks @jim_thompson5910 ;p)

OpenStudy (anonymous):

\[y''=\frac{ -12x }{ (x^2 +1)^2 }\]

OpenStudy (anonymous):

@Yttrium did what?

OpenStudy (yttrium):

Nevermind lol

OpenStudy (anonymous):

Now I don't know what to do. @freckles @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I don't agree with your y '' @Ephemera you're close though

jimthompson5910 (jim_thompson5910):

you should find that \[\Large y \ ' = \frac{6}{1+2x+x^2}\]

jimthompson5910 (jim_thompson5910):

now use the quotient rule

OpenStudy (anonymous):

That's what I did.

OpenStudy (yttrium):

Perhaps you can also use \[y' = 6(x+1)^{-2}\] and simply use power rule.

OpenStudy (anonymous):

Found my mistake

OpenStudy (anonymous):

I did it with (x+1)^2 at the bottom.

OpenStudy (anonymous):

Instead of that I put the original equation

OpenStudy (anonymous):

And got\[\frac{ -12 }{ (x+1)^3 }\]

jimthompson5910 (jim_thompson5910):

oh I see now, I didn't simplify fully yes i'm getting \[\Large y \ '' = \frac{-12}{(x+1)^3}\] as well

OpenStudy (anonymous):

So what is needed from me now?

jimthompson5910 (jim_thompson5910):

it asks where the function is concave up, basically it's asking for what x values makes y '' > 0 true

jimthompson5910 (jim_thompson5910):

the numerator is fixed to being negative, so just focus on the denominator

OpenStudy (anonymous):

\[x < -1\]

OpenStudy (anonymous):

-1 will give you indefinite

OpenStudy (anonymous):

or undefined

OpenStudy (anonymous):

While positive on the bottom will give you a negative

jimthompson5910 (jim_thompson5910):

yep, the denominator is negative when x < -1 that makes the whole right side positive

OpenStudy (anonymous):

Thanks

OpenStudy (md152727):

Please close the question @Ephemera